OpenStudy (babynini):

Falling Ball: When an object is allowed to fall freely near the surface of the earth, the gravitational pull is such that the object falls 16 ft in the first second, 48 ft in the next second, 80ft in the next second, and so on. a) Find the total distance a ball falls in 6seconds b) Find a formula for the total distance the ball falls in n seconds

OpenStudy (babynini):

@jtvatsim Whenever you're free! I'll try to work it out on my own until then :)

OpenStudy (jtvatsim):

Sounds good, be back. :)

OpenStudy (babynini):

a_6=496feet

OpenStudy (babynini):

^ (part a)

OpenStudy (babynini):

b) the formula would be a_n=16+(n-1)(32)

OpenStudy (babynini):

Testing it with a_2 the formula seems to be correct :)

OpenStudy (anonymous):

I got the same thing for b, but 576 for a

OpenStudy (babynini):

hmm let's see where I might've messed up a_6=16 +(6-1)(32) correct yeah?

OpenStudy (anonymous):

to find the 6th term, yes. But I think it's asking for the sum of the first 6 terms

OpenStudy (babynini):

oooh, wow. I completely missed that.

OpenStudy (babynini):

so i must find a_4 and a_5

OpenStudy (babynini):

a_4=112, a_5 = 144

OpenStudy (anonymous):

for this one it's easy enough to do that, but there's a formula to find the sum of the first n terms.\[S _{n}=\frac{ n }{ 2 }(a _{1}+a _{n})\]

OpenStudy (babynini):

oh o.0

OpenStudy (anonymous):

Yes those are right

OpenStudy (babynini):

hm but I should probably use that formula you just used because the prof introduced it to us.

OpenStudy (babynini):

a_6=176 correct?

OpenStudy (anonymous):

yes

OpenStudy (babynini):

so we have s_6=(6/3)(16+176) =3(192) =576

OpenStudy (anonymous):

yep

OpenStudy (babynini):

Fantastic! thank you so much :)

OpenStudy (anonymous):

you're welcome!

OpenStudy (babynini):

oh wait! @peachpi the b) asks "TOTAL distance the ball falls at n seconds"

OpenStudy (babynini):

my formula is not for the sum

OpenStudy (babynini):

Right? So the formula would actually be s_n=(n/2)(16+a_n)

OpenStudy (anonymous):

right, and you'd so substitute the formula for the nth term for a_n

OpenStudy (anonymous):

If you need an explicit formula

OpenStudy (babynini):

ok, it just asks for a formula that could generate any nth term.

OpenStudy (babynini):

Er, should I put the formula to find a_n and also the formula for finding the sums?

OpenStudy (anonymous):

so I think what they're asking is for a formula where you would just stick in the n value and get the sum. Sn = 3(16 + a_n) but since we know a_n = 16+(n-1)(32) Sn = 3(16 + 16+(n-1)(32))

OpenStudy (anonymous):

does that make sense? can't really think of a right way to say it

OpenStudy (babynini):

you mean sn=n/6(16+(n-1)(32)) (not ^3)

OpenStudy (babynini):

ahh yes I see what you did.

OpenStudy (anonymous):

oh, yes. n/2

OpenStudy (babynini):

ok :)

OpenStudy (babynini):

thanks, phew. That was close. I almost put the wrong formula xP

OpenStudy (babynini):

Sn = (n/2)(16 + 16+(n-1)(32)) is the final formula?

OpenStudy (anonymous):

yes. I mean you can do some algebra to make it prettier, but it's correct

OpenStudy (babynini):

oh o.0 we should probably make it prettier.. hah

OpenStudy (babynini):

sn=(n/2)(16+512(n-1)) is that still correct? and simplified enough?

OpenStudy (anonymous):

keep going until you get a quadratic, so distribute 512 and combine like terms

OpenStudy (anonymous):

wait no!

OpenStudy (anonymous):